OpenStudy (eratraya):

first time on open study can any one help me Identify the number that does not belong with the other three. Explain your reasoning. negative 10 over 2, negative 13.4, square root 18, 22.7 repeating 7

OpenStudy (phi):

there are different "types" of numbers. I assume you are studying the types. do you remember any of the types ?

OpenStudy (eratraya):

irrational and rational and repeating decimal and terminating decimal 8th grade pre algebra

OpenStudy (phi):

rational (from ratio ) means you can write the number as a fraction irrational means you can't: you can get them from taking roots, for example square roots the only "tricky" part is decimals where you "know" the next decimal (as in repeating decimals), are rational. care to try to answer the question ?

OpenStudy (eratraya):

Order the set of numbers from least to greatest: 2.71 repeating 71, 2 and 3 over 4, square root 5, 5 over 2 2.71 repeating 71, 2 and 3 over 4, square root 5, 5 over 2. square root 5, 5 over 2, 2.71 repeating 71, 2 3 over 4 5 over 2, square root 5, 2.71 repeating 71, 2 3 over 4 2 3 over 4, 2.71 repeating 71, 5 over 2, square root 5

OpenStudy (eratraya):

its multiple question so i think b

OpenStudy (eratraya):

@phi

OpenStudy (eratraya):

so am i correct or wrong

OpenStudy (phi):

5/2 is 2.5 so the first line has 2.7171... versus 2.5. that is not small to big.

OpenStudy (phi):

the last line has a similar problem so it's down to B or C

OpenStudy (eratraya):

c

OpenStudy (phi):

is sqr(5) bigger or smaller than 2.5 ? we could use a calculator or we could square both numbers: sqr(5) * sqr(5) is 5 (by definition) 2.5 * 2.5 is 6.25 or we could square (5/2) * (5/2) = 25/4 = 6 1/4 or 6.25 either way it shows sqr(5) is smaller than 5/2

OpenStudy (eratraya):

so is it b

OpenStudy (eratraya):

@phi

OpenStudy (eratraya):

can you help

OpenStudy (eratraya):

@.Sam.

OpenStudy (eratraya):

i picked b can you help me with one other